OpenStudy (anonymous):

How do you do this type math problem example 1/5x +8= 3/4 these are fractions if you can't tell.

OpenStudy (mrnood):

\[\frac{ x }{ 5 }+8=\frac{ 3 }{ 4 }\] You need to get x 'on its own' so first multiply the whole equation by 5 to remove the fraction under x write your answer here..

OpenStudy (anonymous):

x/25 + 40= 15/20

OpenStudy (mrnood):

not quite correct - the first term is not right - we are trying to get x on its own

OpenStudy (anonymous):

The teacher showed the class this but I didn't really get it and she couldn't explain it for me understand to help me.

OpenStudy (mrnood):

Well - you have done MOST of the work - I asked you to multiply the equation by 5 (because we have x/5 in the equation) You have made a mistake in the first term: 5 (x/5 + 8) =?

OpenStudy (anonymous):

x/25 + 40 is this part right?

OpenStudy (mrnood):

no - as I said it is the first term 5 (x/5)=?

OpenStudy (anonymous):

The way the teacher said it I thought it went as 1/20x +40

OpenStudy (jdoe0001):

you could just add the fractions bear in mind that \(\bf \cfrac{ x }{ 5 }+8=\cfrac{ 3 }{ 4 }\implies \cfrac{ x }{ 5 }+\cfrac{8}{1}=\cfrac{ 3 }{ 4 }\implies \cfrac{ x }{ 5 }+\cfrac{8}{1}-\cfrac{ 3 }{ 4 }=0\) then just get the LCD and sum the fractions then multiply both sides by the found LCD, then solve for "x"

OpenStudy (anonymous):

I also forgot to add I have to make the problem not a fraction and then solve it

OpenStudy (anonymous):

and thats what I didn't understand

OpenStudy (mrnood):

so why don't you try to do what I asked ? you have nearly done it . I told oyu where the mistake is so try again 5(x/5+8) =?

OpenStudy (anonymous):

is it just x/5 + 40 or x/25 +40 or x/25 +8 because im confused

OpenStudy (jdoe0001):

ohhh... I see...you'd want to get rid of the denominator well... let's look at all 3 denomiators there 5 * 1 * 4 = 20 so let us multiply all by 20 then so we get \(\bf \cfrac{ x }{ 5 }+8=\cfrac{ 3 }{ 4 }\implies \cfrac{ x }{ 5 }+\cfrac{8}{1}=\cfrac{ 3 }{ 4 }\implies {\color{brown}{ \cancel{ 20 }}}\cfrac{ x }{ \cancel{ 5 } }+{\color{brown}{ 20}}\cfrac{8}{1}={\color{brown}{ \cancel{ 20 }}}\cfrac{ 3 }{ \cancel{ 4 } }\implies ?\)

OpenStudy (anonymous):

im 7th grade so this is a little hard to understand

OpenStudy (mrnood):

It's none of yourt previous answers Forget the 8 for a minute what is 5(x/5) (in words that is 5 times x divided by 5)

OpenStudy (jdoe0001):

hmmm at 7th grade... I'd think fractions should have been covered very well by now

OpenStudy (anonymous):

i get distracted by little things easily and the teacher confuses me at some times

OpenStudy (mrnood):

all these excuses are not getting you anywhere - are oyu going to try this - or are oyu waiting for the answer?

OpenStudy (anonymous):

i am trying i really am confused and this is just an example of all the rest of my math hw that i have to do so im also in a hurry.

OpenStudy (jdoe0001):

FallinLeaf may well be, unfortunately, that still doesn't exempt you from studying the material, or try many great tutorials online

OpenStudy (anonymous):

im am trying

OpenStudy (mrnood):

you cannot hurry something that oyu don't know how to do in the original equation ther is x/5 so - to get rid of that we multiply the whole equation by 5 you NEARLY did it in your FIRST reply go back and do it again you were nearly correct

OpenStudy (jdoe0001):

\(\bf \bf \cfrac{ x }{ 5 }+8=\cfrac{ 3 }{ 4 }\implies \cfrac{ x }{ 5 }+\cfrac{8}{1}=\cfrac{ 3 }{ 4 }\implies {\color{brown}{ \cancel{ 20 }}}\cfrac{ x }{ \cancel{ 5 } }+{\color{brown}{ 20}}\cfrac{8}{1}={\color{brown}{ \cancel{ 20 }}}\cfrac{ 3 }{ \cancel{ 4 } }\implies ?\) would give you a plain linear equation with no denominators

OpenStudy (anonymous):

i have good grades but i just don't get this for some reason, and im not good at fractions no matter how hard i try

OpenStudy (mrnood):

so what do you want?

OpenStudy (mrnood):

I have asked oyu 3 times to do something to correct your mistske, but all oyuve done is complain how do you want to solve this?

OpenStudy (anonymous):

im trying ok and ive been trying to answer you ok is it x/5 +40?

OpenStudy (anonymous):

4/20 + 40/20 ? this is just how i think the teacher wanted it so im just asking

OpenStudy (mrnood):

and I already told you that it is not that there are 3 terms in the equation: x/5 8 15/20 multiply each of them by 5 and write the answer here x/5 *5= 8 *5= 15/20 *5 =

OpenStudy (mrnood):

sorry - my error - the three terms are below: multiply each of them by 5 and write the answer here x/5 *5= 8 *5= 3/4 *5 =

OpenStudy (anonymous):

5/25 40 15/20

OpenStudy (mrnood):

when you multiply a fraction by a number you only multiply the TOP number I'm sorry - you seem to be a very long way from being ready to do this type of question. I suggest you revise your fractions multiplying and using LCD to combine fractions. You need to be able to do these things before you can do the whole of this question. I will give you the step I was leading to: 5(x/5+8) = x+40 5(3/4) = 15/4 Please look at your text or ask your teacher again - you need some background study before you are ready to complete this homework

OpenStudy (mrnood):

Giving you the answer is no use to you at all If you want to learn - you have to study and DO the work Do what oyu have to - I will not give you the answer

OpenStudy (anonymous):

sorry but im just tired and i will do what you said thank you a lot

OpenStudy (anonymous):

bye i have to go